OpenStudy (superhelp101):

Factor 4z^2-4z+1 I know that (a+b)^2=a2+2ab+b2 and I know that the answer is (2z-1)^2. I just can't understand how to factor it.

OpenStudy (superhelp101):

@tkhunny

OpenStudy (tkhunny):

For me, I would give it the suspicion test. 4z^2 is a perfect square +1 is a perfect square. The very first thing I would try is: (2z + 1)^2 Does that work? Oh, wait. The second thing I would try, noting the negative middle term, is (2z-1)^2 Does that work? You can ALWAYS use the quadratic formula.

OpenStudy (superhelp101):

I could use (a+b)^2=a2+2ab+b2 right?

OpenStudy (superhelp101):

that's what my teacher wants me to use @tkhunny

OpenStudy (tkhunny):

...or (a-b)^2 = a^2 - 2ab + b^2

OpenStudy (superhelp101):

oh yeah oops, how would I use it though ?

OpenStudy (tkhunny):

Essentially what I did in the Suspicion Method. Establish the correct pattern and apply the principle.

OpenStudy (superhelp101):

So a=4z_ and b=1_

OpenStudy (superhelp101):

@tkhunny

OpenStudy (tkhunny):

That's what I did first, too. Now, figure out that we need a negative and change b.

OpenStudy (superhelp101):

I got (4z-1)^2 when I plug in

OpenStudy (superhelp101):

I don't understand how I can get from their to (2z-1)^2 . Do you mind teaching me? @tkhunny

OpenStudy (tkhunny):

Seriously, you just have to recognize the pattern. (a-b)^2 = a^2 - 2ab + b^2 and a^2 - 2ab + b^2 = (a-b)^2 It works both ways.

OpenStudy (superhelp101):

but let's say they didn't give me the answer: (2z-1)^2. How would I do it?

OpenStudy (tkhunny):

We have been over this. My suspicion method IS how I would do it. Recognize the pattern and implement the relationship. Alternatively, use the quadratic formula. If you were to solve, 4z^2-4z+1 = 0, how would you do it?

OpenStudy (xapproachesinfinity):

why don't you treat 2z as big X then it is clear that \(X^2-2X+1=(X-1)^2\)

OpenStudy (superhelp101):

my teacher tells me not to use the quadratic formula @tkhunny @xapproachesinfinity but really I don't have 2z. I got the answer cuz my teacher when over it but I am suppose to solve it without knowing that it equals (2z-1)^2

OpenStudy (tkhunny):

i find it particularly odd to tell you not to use tools you are taught. Substitution is a very effective method. Try it.

OpenStudy (xapproachesinfinity):

well you have to look for 2z there you have 4z^2=(2z)^2 you need to look for perfect square as @tkhunny said and 1 is already good

OpenStudy (tkhunny):

Essentially, what you are asking is how to solve the problem without using any known method. You need to see that this is an impossible assignment. Recognize the form. Make a substitution. Use the quadratic formula. Ask a friend. So far, everything you suggest it immediately ruled out. You must do SOMETHING or you cannot solve the problem. There is not a method called "Fall Out of the Sky". That doesn't work at all.

OpenStudy (xapproachesinfinity):

see this example: 9c^2+12c+4

OpenStudy (xapproachesinfinity):

how would you think about this?

OpenStudy (xapproachesinfinity):

i would have to agree with @tkhunny , you must use one way or another to solve this otherwise it is impossible

OpenStudy (superhelp101):

that's making me feel dumb @tkhunny lol ok so for that example I would use (a+b)^2=a2+2ab+b2 am I right?

OpenStudy (xapproachesinfinity):

right! that the essential of the essentials to recognize perfect square actually it is not essential is the heart of this haha

OpenStudy (tkhunny):

No need to feel dumb. That is a voluntary feeling. Just don't do it. There is room in this world for just "recognize" the pattern.

OpenStudy (xapproachesinfinity):

^_^! if you don't get stuck you don't learn!

OpenStudy (superhelp101):

haha @xapproachesinfinity :)

OpenStudy (superhelp101):

What would i do next?

OpenStudy (xapproachesinfinity):

now you need to recognize two perfect squares

OpenStudy (xapproachesinfinity):

9c^2=? 4=?

OpenStudy (superhelp101):

3c and 2

OpenStudy (xapproachesinfinity):

correct! so we could have (3c+2)^2 or (3c-2)^2

OpenStudy (xapproachesinfinity):

what you need to check is if the middle term is 2ab

OpenStudy (superhelp101):

12c

OpenStudy (xapproachesinfinity):

so it is good! now we have +12c which one these (3c+2)^2 , (3c-2)^2 would give us +12c in the middle recall (a+b)^2=a^2+2ab+b^2 (a-b)^2=a^2-2ab+b^2

OpenStudy (superhelp101):

(3c+2)^2

OpenStudy (xapproachesinfinity):

correct!

OpenStudy (superhelp101):

yayy! thanks you so much! :D

OpenStudy (xapproachesinfinity):

yw! with pratice you went need to go over all of this right away you see what you need to do! it is brain training lol

OpenStudy (xapproachesinfinity):

practice*

OpenStudy (xapproachesinfinity):

won't*

OpenStudy (superhelp101):

yes so true!! lol :)

OpenStudy (xapproachesinfinity):

darn it! i'm pretty bad a typing haha